# The Dynamic Net Architecture: Learning Robust and Holistic Visual Representations Through Self-Organizing Networks

### v1

@@ Appendix E Alternative Neuron Selection
Only one of a set of alternative neurons is allowed to be active. The selection process is determined by evaluating the correlation between feature patches and alternative kernel filters. Initially, patches matching the size of the feature kernel are extracted from every position in the input $(\boldsymbol{y}^{(S1)};\boldsymbol{y}^{(S2)}[t-1])$. Subsequently, the correlation of these patches with the feature kernels is computed, and at each location, the alternative kernel exhibiting the highest correlation is chosen. Consequently, multiple features may be active at each location, but only one of the competing feature channels — the one best matching a given input patch — achieves activation. At initialization (see Appendix B), where all alternative channels fit equally well to a given local pattern, the first channel is selected. However, after applying a Hebbian update, this channel starts specializing in recognizing a specific local pattern. Consequently, the activation of subsequent local patterns may either trigger the same channel if they have a high correlation or activate another channel better suited to recognizing the new pattern.
@@ Appendix F Feature Extraction
The input images have $C=1$ channel and $C^{(S1)}=4$ features are extracted at each spatial location. We utilize hand-crafted filters with a size of $5\times 5$ as displayed in Figure 9, each of the four filters corresponding to a particular line orientation (vertical, positive diagonal, horizontal, and negative diagonal). This approach is motivated by the observation that, given our specific elementary feature types (cf. Fig. 9), straight lines, when viewed locally, can be expressed as combinations of these fundamental line types (e.g., a line with a $20^{\circ}$ angle activates the horizontal and the $+45^{\circ}$ diagonal feature neurons). The use of such manually designed filters allows a more straightforward interpretation of the results. While more complex datasets necessitate learned filters, our primary focus in this work is on constructing net fragments, and thus, we favor the simplicity of hand-crafted filters.
These filters move across the image with a step size (“stride”) of $1$ (assuming 00 as input signal beyond the image border: “zero padding”) so that the input image and output image are of the same size. Thus, the input $\boldsymbol{x}$ in the feature extraction stage are images of size $(1\times 32\times 32)$ and the output $\boldsymbol{y}^{(S1)}$ feature activation maps of size $(4\times 32\times 32)$. The dataset’s lines activate multiple filters at different positions, facilitating the construction of net fragments across these feature channels. We convert the floating-point output of the convolutional weights to binary activations using binary neurons $\mathrm{B}()$ with a bias of $b^{(S1)}=0.5$, i.e. neurons fire a binary spike if $\boldsymbol{a}_{j}>0.5$ (see eq. 2).
@@ Appendix G Parameters
We train the model with a learning rate of $\alpha=0.2$ for $100$ training cycles (epochs). The model uses $n_{a}=10$ alternative neurons with lateral connections spanning a kernel of $11\times 1$ 11 neurons. Since we use $C^{(S1)}=4$ filters for feature extraction in $S1$ (see Appendix F), the number of alternative channels in $S2$ corresponds to $C^{(S2)}=4\cdot n_{a}=40$. The resulting weight matrices have dimensions $\boldsymbol{W}^{(F)}\in\mathbb{R}^{40\times 4\times 11\times 11}$ and $\boldsymbol{W}^{(L)}\in\mathbb{R}^{40\times 40\times 11\times 11}$, respectively $\boldsymbol{W}^{(S2)}\in\mathbb{R}^{40\times 44\times 11\times 11}$.
Appendix H Learned Support Strength
Figure 10: The average support active and inactive neurons receive during training, captured by $\boldsymbol{a}^{\prime}_{j}$. The support strength is shown on the y-axis, and the training progress (epoch) is on the x-axis.
A crucial parameter for gauging the efficacy of net fragments involves quantifying the difference in support received by active and inactive neurons. During the initial stages of training, active neurons receive a support of $1$, and inactive neurons receive a support of 00. The support increases during training as net fragments are formed. Figure 10 visually represents the average support for active and inactive neurons, captured by $\boldsymbol{a}^{\prime}_{j}$ (see eqs. 3, 5). Notably, the graph demonstrates a widening gap between the support for active and inactive neurons. This discrepancy suggests an enhancement in the robustness of statistically relevant patterns over time as active neurons receive increased support and the support of inactive neurons remains close to 00.
@@ Appendix I Quantifying Noise Filtration
Figure 11: System performance as function of the noise level. First row: noise reduction rate; second row: recall; third row: precision. Columns differ in the inhibition coefficient $\gamma$ that models inhibition. The colors within each plot represent different activation biases $b^{(S2)}$.
@@
Rconstructing Subtractive Noise.
Figure 12: Reconstruction of partial occlusion. First row: feature reconstruction rate; second row: recall; third row: precision. Columns are distinguished by the inhibition coefficient $\gamma$, and line colors within plots distinguish activation bias $b^{(S2)}$.
@@
In the case of the depicted straight lines, the feature neuron activations of the two most similar lines overlap with $21.2\%$. Therefore, achieving precision and recall values surpassing this threshold permits reliable discrimination even among similar lines. Hence, we posit that various of the tested parameter configurations are well suited as they permit this distinction.

### v2

@@ Appendix D Alternative Neuron Selection
Only one of a set of alternative neurons is allowed to be active. The selection process is determined by evaluating the correlation between feature patches and alternative kernel filters. Initially, patches matching the size of the feature kernel are extracted from every position in the input $(\boldsymbol{y}^{(S1)};\boldsymbol{y}^{(S2)}[t-1])$. Subsequently, the correlation of these patches with the feature kernels is computed, and at each location, the alternative kernel exhibiting the highest correlation is chosen. Consequently, multiple features may be active at each location, but only one of the competing feature channels — the one best matching a given input patch — achieves activation. At initialization (see Appendix A), where all alternative channels fit equally well to a given local pattern, the first channel is selected. However, after applying a Hebbian update, this channel starts specializing in recognizing a specific local pattern. Consequently, the activation of subsequent local patterns may either trigger the same channel if they have a high correlation or activate another channel better suited to recognizing the new pattern.
Appendix E Dataset
The images are generated when required and may vary from one epoch to another. During each training cycle, we randomly generate $300$ image instances (each line about $5$ times). During evaluation, we sample each of the $59$ distinct lines exactly once. In contrast to training, the samples used for evaluation comprise local distortions in the form of additive Gaussian nose and missing line segments (subtractive noise), and we evaluate the net fragments’ capability to remove these.
To introduce additive Gaussian noise to the afferent input in stage $S2$, we probabilistically flip neurons within each channel with a probability of up to $20\%$. Since we use four feature channels in our experiments, this corresponds to a probability of $1-(1-0.2)^{4}=59\%$ that a neuron is flipped at any spatial location. To evaluate subtractive noise, we create discontinuous lines by deleting a line segment of $1$ to $7$ pixels in the middle.
To demonstrate the generalization capability of CNA areas, we generate validation samples that differ from the ones used during training. Specifically, we produce binary images measuring ($32\times 32$) pixels depicting kinked lines by generating four random points within the image and drawing straight lines between these points.
Additionally, we generate binary images measuring ($64\times 64$) pixels by hand that represent more complex structures. These images include line drawings of digits 00-$9$, characters, and objects such as a house, a star, a wheel, a clock, etc. Some of these images are shown in Figure 4, and an extensive list can be found at https://github.com/sagerpascal/cna.
@@ Appendix F Feature Extraction
The input images have $C=1$ channel and $C^{(S1)}=4$ features are extracted at each spatial location. We utilize hand-crafted filters with a size of $5\times 5$ as displayed in Figure 10, each of the four filters corresponding to a particular line orientation (vertical, positive diagonal, horizontal, and negative diagonal). This approach is motivated by the observation that, given our specific elementary feature types (cf. Figure 10), straight lines, when viewed locally, can be expressed as combinations of these fundamental line types (e.g., a line with a $20^{\circ}$ angle activates the horizontal and the $+45^{\circ}$ diagonal feature neurons). The use of such manually designed filters allows a more straightforward interpretation of the results. While more complex datasets necessitate learned filters, our primary focus in this work is on constructing net fragments, and thus, we favor the simplicity of hand-crafted filters.
These filters move across the image with a step size (‘stride’) of $1$ (assuming 00 as input signal beyond the image border: ‘zero padding’) so that the input image and output image are of the same size. Thus, the input $\boldsymbol{x}$ in the feature extraction stage are images of size $(1\times 32\times 32)$ and the output $\boldsymbol{y}^{(S1)}$ feature activation maps of size $(4\times 32\times 32)$. The dataset’s lines activate multiple filters at different positions, facilitating the construction of net fragments across these feature channels. We convert the floating-point output of the convolutional weights to binary activations using binary neurons $\mathrm{B}()$ with a bias of $b^{(S1)}=0.5$, i.e. neurons fire a binary spike if $\boldsymbol{a}_{j}>0.5$ (see eq. 2).
@@ Appendix G Parameters
We train the model with a learning rate of $\alpha=0.2$ for $100$ training cycles (epochs). The model uses $n_{a}=10$ alternative neurons with recurrent connections spanning a kernel of $11\times 1$ 11 neurons. Since we use $C^{(S1)}=4$ filters for feature extraction in $S1$ (see Appendix F), the number of alternative channels in $S2$ corresponds to $C^{(S2)}=4\cdot n_{a}=40$. The resulting weight matrices have dimensions $\boldsymbol{W}^{(F)}\in\mathbb{R}^{40\times 4\times 11\times 11}$ and $\boldsymbol{W}^{(L)}\in\mathbb{R}^{40\times 40\times 11\times 11}$, respectively $\boldsymbol{W}^{(S2)}\in\mathbb{R}^{40\times 44\times 11\times 11}$.
Appendix H Autoencoder Parameter
We compare our CNA model to an ANN, specifically an autoencoder (Hinton and Salakhutdinov, 2006), which is a widely used self-supervised technique for generating input representations. An autoencoder is particularly suited for comparison because it includes a built-in decoder, allowing for feature reconstruction and visualization, which is required for feature comparison. To ensure a fair comparison, we also employ convolutional filters (LeCun et al., 1989) to the autoencoder and incorporate state-of-the-art design principles (we even permit more data and feature channels to the autoencoder).
The autoencoder is trained to reconstruct the feature activations, $\boldsymbol{y}^{(S1)}$, using the same training dataset as our CNA model. It consists of an 8-layer architecture, with an encoder composed of convolutional layers with a kernel size of $3$ and a stride of $2$, with progressively increasing channel counts of $32$, $64$, $128$, and $256$ layers. The decoder mirrors this structure, using transposed convolutional layers with $256$, $128$, $64$, and $32$ channels, respectively. ReLU activations are applied after each layer. Consequently, the autoencoder is a relatively large model with $777$k parameters, utilizing up to $256$ channels, whereas the CNA model operates with only $44$ channels.
The autoencoder is trained by minimizing the L2 norm between the input features and their reconstructed counterparts, using the Adam optimizer (Kingma and Ba, 2015) with a learning rate of $1\times 10^{-4}$, beta parameters of $(0.9,0.999)$, and a weight decay of $1\times 10^{-8}$.
To provide a robust comparison, we increase the number of training samples for the autoencoder. Rather than training for 100 epochs with 300 samples per epoch as done for the CNA model, we train the autoencoder for $200$ epochs, using $10,000$ images per epoch. This results in a significantly larger training set but ensures that the autoencoder achieves a highly accurate reconstruction. The autoencoder is trained with a batch size of $512$.
@@ Appendix I Quantifying Noise Filtration
Figure 11: System performance as function of the noise level. First row: noise reduction rate; second row: recall; third row: precision. Columns differ in the inhibition coefficient $\gamma$ that models inhibition. The colors within each plot represent different activation biases $b^{(S2)}$. The dotted red line is the autoencoder, serving as baseline.
@@
Furthermore, all CNA models (except the one with the lowest power factor and bias) outperform the autoencoder, demonstrating that they achieve performance that is competitive with state-of-the-art deep learning models.
Reconstructing Subtractive Noise.
Figure 12: Reconstruction of partial occlusion. First row: feature reconstruction rate; second row: recall; third row: precision. Columns are distinguished by the inhibition coefficient $\gamma$, and line colors within plots distinguish activation bias $b^{(S2)}$. The dotted red line is the autoencoder model.
@@
Compared to the autoencoder baseline, the CNA model achieves higher recall and precision but a lower feature reconstruction rate. This indicates that the autoencoder is superior at reconstructing the pixels that have been removed but that, in general, much more noise (other neurons are activated/deactivated) is introduced.
@@
In the case of the depicted straight lines, the feature neuron activations of the two most similar lines overlap with $21.2\%$. Therefore, achieving precision and recall values surpassing this threshold permits reliable discrimination even among similar lines, which is the case for the CNA but not for the autoencoder (e.g., see Figure 11).